OpenStudy (anonymous):

A family has two cars. The first car has a fuel efficiency of 40 miles per gallon of gas and the second has a fuel efficiency of 15 miles per gallon of gas. During one particular week, the two cars went a combined total of 1900 miles, for a total gas consumption of 60 gallons. How many gallons were consumed by each of the two cars that week? Firstcar:gallons Secondcar:gallons

OpenStudy (aveline):

Can you write two equations to model the situation?

OpenStudy (jdcap):

\[40f+15s=1900 \hspace{0.5cm}Eq.1\\ f+s=60 \hspace{1cm}Eq.2 \\ \color{orange}{Eq.1-40\times Eq.2}\\-25s=-500\Rightarrow \,\color{blue}{s=20}\\f=60-20\Rightarrow \color{blue}{f=40}\]